OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - x^2-1 x-1 -------- / ----- 2x 4x^2 **Note the / means to divide

satellite73 (satellite73):

\[\frac{x^2-1}{2x}\times \frac{4x^2}{x-1}\] is a start then factor and cancel

OpenStudy (abbycross167):

Could you show me how to do that ?

satellite73 (satellite73):

\[x^2-1=(x+1)(x-1)\] i guess is the only thing you can factor more cancels though

satellite73 (satellite73):

\[\frac{(x+1)(x-1)}{2x}\times \frac{4x^2}{x-1}\]

satellite73 (satellite73):

the \(x-1\) cancels top and bottom as does the \(\frac{4x^2}{2x}\) which becomes \(2x\)in the numerator

OpenStudy (abbycross167):

Thank you for the help!!! I really appreciate it!!